### v2

# Parallel transport in rotating frames and projective holonomic quantum computation

Abstract
Nonadiabatic holonomic quantum computation is a promising approach for implementing quantum gates, offering both efficiency and robustness against certain types of errors. A key element of this approach is a geometric constraint known as the parallel transport condition. In line with the principle of covariance, this condition must be appropriately modified when changing reference frames. In this paper, we detail how to adjust the parallel transport condition when transitioning from the laboratory frame to a rotating reference frame. Furthermore, building on gauge invariance considerations, we develop a framework for nonadiabatic holonomic quantum computation with projective gates. The parallel transport condition of this framework effectively addresses the issue of global dynamical phases inherent in conventional nonadiabatic holonomic quantum computation. Finally, we extend the isoholonomic inequality to projective gates, establishing a minimum execution time for projective holonomic quantum gates.
Holonomic quantum computation; parallel transport; holonomic gate; nonabelian geometric phase; isoholonomic inequality; isoholonomic problem; quantum speed limit.
I   Introduction
Nonadiabatic holonomic quantum computation [1, 2, 3] has been introduced as a fast alternative to its well-established, robust, but relatively slow adiabatic counterpart [4, 5, 6]. While adiabatic holonomic quantum computation relies on the gradual manipulation of control parameters, nonadiabatic holonomic quantum computation implements quantum gates through internal dynamics driven by parallel transporting Hamiltonians. These Hamiltonians, fundamental to the geometric nature of the gates, evolve the computational subspace without inducing any time-local rotations within it, ensuring that the computation is governed purely by geometric phases rather than dynamical effects.
In the laboratory frame, the parallel transport condition can be formulated as the requirement that the velocity of any computational basis remains within the kernel bundle of the non-Abelian Aharonov-Anandan connection [7, 8]. To simplify the analysis of quantum dynamical equations, a common approach is to switch to a rotating reference frame. However, to maintain physical consistency, the dynamical equations must be transformed covariantly when changing reference frames. After outlining key concepts in holonomic quantum computing, we work out how to transform the parallel transport condition when transitioning from the laboratory frame to a rotating reference frame.
Rotation operators in the center of the unitary group leave quantum states and observables invariant due to the global $\mathrm{U}(1)$ symmetry of quantum systems. Requiring that the parallel transport condition remains invariant under such central rotations naturally leads to a projective version of the parallel transport condition. This projective condition effectively resolves the problem of global dynamical phases in holonomic quantum computation and has been widely, though often implicitly, applied in practice [10, 11, 12, 13, 9, 14, 15, 16, 17]. We explicitly derive the projective parallel transport condition from a gauge-theoretic framework for holonomic quantum computation with projective gates.
The isoholonomic inequality, a fundamental result in holonomic quantum computation [18, 19, 20], establishes a sharp lower bound on the length of the loops along which the computational subspace can be parallel transported to holonomically implement a given quantum gate [20]. We extend this inequality to projective quantum gates and, as a corollary, provide an estimate for the execution time of projective holonomic quantum gates. This extension offers deeper insight into the geometric constraints governing the implementation of quantum gates in the projective setting.
Remark.
In this paper we use the word frame in two different senses. First, there are frames of reference used to describe the entire quantum system, referred to as laboratory or rotating frames. These provide a basis for understanding the system’s overall behavior. Second, we introduce frames that span the computational space, which we call computational frames or $n$-frames. The context should make it clear which type of frame is being discussed.
II   Conventional nonadiabatic holonomic quantum computation
Holonomic quantum computing uses holonomy to implement gates. Holonomy refers to the geometric phenomenon that when the computational space is transported along a closed path, a co-transported computational frame, that is, a frame spanning the computational space, may end up rotated relative to its original configuration, even though the transport is parallel, that is, performed without causing any time-local rotation of the computational frame.
A holonomic gate is a unitary operator representing the rotation of a computational frame resulting from a cyclic parallel transport of the computational space. A more precise description of a holonomic gate is as follows (see also [1, 2, 3, 20]). Consider a quantum system with a Hilbert space $\mathcal{H}$ and a computational subspace $\mathcal{R}$ [^1] Thus, computational input states are prepared to have support in $\mathcal{R}$, and are manipulated by unitary operators on $\mathcal{R}$, a.k.a gates. Let $\Pi_{t}$ be a one-parameter family of isometric embeddings of $\mathcal{R}$ in $\mathcal{H}$ that transport $\mathcal{R}$ in a loop:
$\mathcal{R}_{t}=\Pi_{t}(\mathcal{R}),\qquad\mathcal{R}_{0}=\mathcal{R}_{\tau}=\mathcal{R}.$ (1)
The transport is said to be parallel if, for each vector $|\psi\rangle$ in $\mathcal{R}$ and at every time $t$, the curve $|\psi_{t}\rangle=\Pi_{t}|\psi\rangle$ intersects the displaced computational space perpendicularly:
[…]
The final operator $\Pi_{\tau}$, which maps $\mathcal{R}$ isometrically onto itself, is the holonomic gate associated with the loop $\mathcal{R}_{t}$.
Geometrically, condition (2) means that the operators $\Pi_{t}$ do not cause any time-local rotation within the computational space. The resulting rotation of a computational frame is thus solely a consequence of the translational motion of $\mathcal{R}$ through the Hilbert space. A crucial fact is that any curve of equidimensional subspaces of $\mathcal{H}$ can be realized as a transport by a unique one-parameter family of parallel transporting operators. These operators are the parallel transport operators associated with the curve. [^2] We assume that the parameter of any one-parameter family—a curve—of states, operators, or spaces ranges from 00 to $\tau$, and that the family depends piecewise smoothly on the parameter.
In practice, nonadiabatic holonomic gates are implemented by configuring the system’s Hamiltonian such that the time evolution operator parallel transports the computational space along a closed path, resulting in the desired gate as the holonomy. A Hamiltonian $H_{t}$ is said to be parallel transporting if its associated time evolution operator $U_{t}$ displaces the computational space in a parallel manner. An equivalent condition is that for any orthonormal basis $|v_{1}\rangle,|v_{2}\rangle,\dots,|v_{n}\rangle$ in the computational space,
$\langle v_{k}|U_{t}^{\dagger}H_{t}U_{t}|v_{l}\rangle=0\quad(k,l=1,2\dots,n)$ (3)
holds at all times [1, 2, 3, 20]. The parallel transport condition (3) ensures the Hamiltonian induces no couplings within the computational space.
II.A   Gauge theoretic description of conventional nonadiabatic holonomic computation
An $n$-frame is a sequence of $n$ orthonormal vectors in $\mathcal{H}$. The set of all $n$-frames in $\mathcal{H}$ forms the Stiefel manifold $\mathcal{V}(n;\mathcal{H})$. In this paper, we will represent $n$-frames as row matrices, $V=(|v_{1}\rangle\,|v_{2}\rangle\dots|v_{n}\rangle)$. This allows us to apply standard matrix algebra to $n$-frames. In particular, we can act component-wise on an $n$-frame from the left with an operator, thereby generating a new row matrix of vectors that may or may not be a new $n$-frame. [^3] The resulting matrix of vectors is an $n$-frame if and only if the operator is an isometry on the span of the input frame. We can also act on an $n$-frame from the right with a real or complex matrix (with $n$ rows) and generate a new row matrix of vectors that are linear combinations of the original frame’s components; see [20] for details.
The Grassmann manifold $\mathcal{G}(n;\mathcal{H})$ is the manifold of $n$-dimensional subspaces of $\mathcal{H}$. Its topology and smooth structure are derived from its identification with the space of orthogonal projection operators on $\mathcal{H}$ of rank $n$, where each subspace is identified with the orthogonal projection onto that subspace.
Suppose the computational space $\mathcal{R}$ has dimension $n$, making it an element of $\mathcal{G}(n;\mathcal{H})$. A smooth transformation of $\mathcal{R}$ corresponds to a curve $\mathcal{R}_{t}$ in $\mathcal{G}(n;\mathcal{H})$, or equivalently to a curve of orthogonal projection operators $P_{t}$ of rank $n$, where $P_{t}$ is the orthogonal projection onto $\mathcal{R}_{t}$. The transformation is said to be closed if $\mathcal{R}_{\tau}=\mathcal{R}$.
The linear span of each $n$-frame is an element of $\mathcal{G}(n;\mathcal{H})$, and the surjective map
$\mathcal{V}(n;\mathcal{H})\ni V\mapsto\operatorname{span}V\in\mathcal{G}(n;\mathcal{H})$ (4)
is a principal bundle called the Stiefel-Grassmann bundle. The symmetry group of this bundle is the group of unitary $n\times n$ matrices $\mathrm{U}(n)$, which acts from the right on $\mathcal{V}(n;\mathcal{H})$.
Let $\mathfrak{u}(n)$ be the Lie algebra of $\mathrm{U}(n)$, consisting of the skew-Hermitian $n\times n$ matrices. The $\mathfrak{u}(n)$-valued Aharonov-Anandan connection $\mathcal{A}$ on $\mathcal{V}(n;\mathcal{H})$ is defined as
$\mathcal{A}(X)=V^{\dagger}X\quad\big{(}X\in\mathrm{T}_{V}\mathcal{V}(n;\mathcal{H})\big{)}.$ (5)
We say that a tangent vector of $\mathcal{V}(n;\mathcal{H})$ is horizontal if it is annihilated by $\mathcal{A}$, and we say that a curve in $\mathcal{V}(n;\mathcal{H})$ is horizontal if all its velocity vectors are horizontal. A fundamental result from the theory of fiber bundles asserts that for every smooth curve $\mathcal{R}_{t}$ in $\mathcal{G}(n;\mathcal{H})$ and every $n$-frame $V_{0}$ spanning $\mathcal{R}_{0}$, there is a unique horizontal curve $V_{t}$ in $\mathcal{V}(n;\mathcal{H})$ that starts at $V_{0}$ and is such that $V_{t}$ spans $\mathcal{R}_{t}$; see [21]. This curve is the horizontal lift of $\mathcal{R}_{t}$ starting at $V_{0}$. If $\mathcal{R}_{t}$ is a closed curve at $\mathcal{R}$, and $V_{t}$ is its horizontal lift, then both $V_{0}$ and $V_{\tau}$ span $\mathcal{R}$. The operators $\Pi_{t}=V_{t}V_{0}^{\dagger}$ are the parallel transport operators associated with $\mathcal{R}_{t}$, and the final operator $\Pi_{\tau}=V_{\tau}V_{0}^{\dagger}$ is the holonomy.
II.B   The isoholonomic inequality and an execution time estimate for holonomic gates
We define the length of a curve $\mathcal{R}_{t}$ in $\mathcal{G}(n;\mathcal{H})$ as
$\operatorname{Length}[\mathcal{R}_{t}]=\int_{0}^{\tau}\mathrm{d}t\,\sqrt{\tfrac{1}{2}\operatorname{tr}\big{(}\dot{P}_{t}^{2}\big{)}},$ (6)
where $P_{t}$ is the orthogonal projection onto $\mathcal{R}_{t}$. Suppose $\mathcal{R}_{t}$ is closed with holonomy $\Gamma$. The isoholonomic inequality for quantum gates states that the length of $\mathcal{R}_{t}$ is always greater than or equal to the isoholonomic bound
$L(\Gamma)=\sqrt{\sum_{j=1}^{n}\theta_{j}(2\pi-\theta_{j})}.$ (7)
Here, $\theta_{1},\theta_{2},\dots,\theta_{n}$ are the phases of the eigenvalues of $\Gamma$ in the interval $[0,2\pi)$; see [20] for a derivation of this fact.
From the isoholonomic inequality it follows that if a Hamiltonian $H_{t}$ generates the curve $\mathcal{R}_{t}$, meaning that $\mathcal{R}_{t}=U_{t}(\mathcal{R})$ where $U_{t}$ is the time-evolution operator associated with $H_{t}$, then the evolution time $\tau$ is not less than
$\tau[H_{t};\Gamma]=\frac{L(\Gamma)}{\langle\!\langle\sqrt{I(H_{t};\mathcal{R}_{t})}\rangle\!\rangle}.$ (8)
The denominator is the time average of the square root of the skewness measure
$I(H_{t};\mathcal{R}_{t})=-\frac{1}{2}\operatorname{tr}\big{(}[H_{t},P_{t}]^{2}\big{)}$ (9)
[…]
II.C   Parallel transport in a rotating frame
Equation (3) describes the parallel transport condition relative to the laboratory frame. If the reference frame is changed, the parallel transport condition must be modified covariantly to maintain physical consistency. Here, we derive the equation that must be satisfied in a rotating reference frame to ensure that the parallel transport condition is satisfied in the laboratory frame.
Let $R_{t}$ be a family of unitaries varying smoothly with $t$. Consider the system in the rotating frame picture specified by the following transformation rules for state vectors, observables, and the Hamiltonian:
$|\psi_{t}^{\textsc{rf}}\rangle$ $=R_{t}|\psi_{t}\rangle,$ (10)
$O^{\textsc{rf}}_{t}$ $=R_{t}O_{t}R^{\dagger}_{t},$ (11)
$H^{\textsc{rf}}_{t}$ $=R_{t}H_{t}R^{\dagger}_{t}+i\dot{R}_{t}R_{t}^{\dagger}.$ (12)
The potential $A_{t}=i\dot{R}_{t}R_{t}^{\dagger}$ compensates for the additional dynamics introduced by the time-varying reference frame.
Let $V$ be any computational frame, and write $|v_{j;t}\rangle$ for the evolution of the $j$th component of $V$. Then
$\langle v_{k;t}|\dot{v}_{l;t}\rangle=\langle v_{k;t}^{\textsc{rf}}|d_{t}+iA_{t}|v_{l;t}^{\textsc{rf}}\rangle,$ (13)
[…]
$\langle v_{k;t}^{\textsc{rf}}|d_{t}+iA_{t}|v_{l;t}^{\textsc{rf}}\rangle=0\quad(k,l=1,2\dots,n).$ (14)
Furthermore, the Hamiltonian satisfies the parallel transport condition in the laboratory frame if and only if
$\langle v_{k;t}^{\textsc{rf}}|H^{\textsc{rf}}_{t}-A_{t}|v_{l;t}^{\textsc{rf}}\rangle=0\quad(k,l=1,2\dots,n).$ (15)
Equation (15) is the equivalent form of the parallel transport condition in the rotating frame picture.
When changing the reference frame, care must also be taken to ensure that the computational space is transported in a loop, because if one of $\mathcal{R}_{t}$ and $\mathcal{R}_{t}^{\textsc{rf}}=R_{t}(\mathcal{R}_{t})$ forms a closed curve in $\mathcal{G}(n,\mathcal{H})$, the other may not necessarily do so. We have that $\mathcal{R}_{t}$ is a closed curve if and only if
$\mathcal{R}_{\tau}^{\textsc{rf}}=R_{\tau}R_{0}^{\dagger}(\mathcal{R}_{0}^{\textsc{rf}}).$ (16)
The computational space being transported in a loop can also be formulated more invariantly: Let $P$ be the orthogonal projection onto $\mathcal{R}$. That the computational space is transported in a loop over time $\tau$ can be expressed in the laboratory frame as $\langle\psi_{\tau}|P|\psi_{\tau}\rangle=1$ for all unit vectors $|\psi_{0}\rangle$ in $\mathcal{R}$. In the rotating frame picture, observables change with time according to Eq. (11), ensuring that all statistical quantities remain invariant under the change of reference frame. Consequently, the computational space is transported in a loop if and only if $\langle\psi_{\tau}^{\textsc{rf}}|P_{\tau}^{\textsc{rf}}|\psi_{\tau}^{\textsc{rf}}\rangle=1$ for all unit vectors $|\psi_{0}^{\textsc{rf}}\rangle$ in $\mathcal{R}_{0}^{\textsc{rf}}$.
II.D   Central rotations as gauge symmetries
A rotational transformation that belongs to the center of the group of unitary operators on $\mathcal{H}$ is a phase factor multiple of the identity operator,
$R_{t}=e^{i\theta_{t}}\mathbbm{1}.$ (17)
Such a transformation does not affect the system’s dynamics because it only adds a global phase to state vectors and thus leave states and observables unchanged. However, it adds a multiple of the identity to the Hamiltonian,
$H_{t}^{\textsc{rf}}=H_{t}+\epsilon_{t}\mathbbm{1}\quad(\epsilon_{t}=-\dot{\theta}_{t}).$ (18)
This means the transformation shifts all energy levels at time $t$ by the value $\epsilon_{t}$. Since only relative phases are detectable, and only differences between energy levels are measurable, this type of transformation can be regarded a gauge symmetry of the system. [^7] This symmetry allows us to choose the zero energy level arbitrarily.
The parallel transport condition (3) is not invariant under such gauge transformations. This is evident from the fact that $H_{t}$ and $H_{t}+\epsilon_{t}\mathbbm{1}$ cannot both satisfy the parallel transport condition in the laboratory frame, despite generating identical dynamics. In the next section, we will derive a condition for parallel transport that remains invariant under the symmetry transformations described by Eq. (17).
III   Projective nonadiabatic holonomic quantum computation
In this section, we derive a projective version of the parallel transport condition from a gauge theory for projective holonomic quantum computation, that is, computation with holonomic gates that are projective unitary operators. The projective parallel transport condition is such that a Hamiltonian $H_{t}$ is parallel transporting if and only if $H_{t}+\epsilon_{t}\mathbbm{1}$ is parallel transporting for every real-valued $\epsilon_{t}$.
III.A   Projective gates
Two gates representing the same projective unitary operator transform states with support in $\mathcal{R}$ identically and are, therefore, computationally interchangeable. We say that the gates represent the same projective gate. Formally, a projective gate on $\mathcal{R}$ is an equivalence class of the relation
$\Gamma_{1}\sim\Gamma_{2}\iff\Gamma_{1}=e^{i\theta}\Gamma_{2}\text{ for some }\theta\in\mathbb{R}$ (19)
on the gates of $\mathcal{R}$. The projective gates form a group that acts on density operators with support in $\mathcal{R}$ as
$\bar{\Gamma}(\rho)=\Gamma\rho\Gamma^{\dagger},$ (20)
where $\Gamma$ is any representative of $\bar{\Gamma}$. Abstractly, the group of projective gates on $\mathcal{R}$ is the holomorphic isometry group on the projective space formed by the unit rank density operators with support in $\mathcal{R}$.
III.B   The projective Stiefel-Grassmann bundle
The center of the symmetry group of the Stiefel-Grassmann bundle consists of all phase factor multiples of the $n\times n$ identity matrix,
$\mathrm{Z}(n)=\big{\{}e^{i\theta}\mathbbm{1}_{n}:\theta\in\mathbb{R}\big{\}}.$ (21)
The center is a Lie subgroup of $\mathrm{U}(n)$ with Lie algebra
$\mathfrak{z}(n)=\{i\epsilon\mathbbm{1}_{n}:\epsilon\in\mathbb{R}\}.$ (22)
The projective Stiefel manifold is the manifold of orbits under the action by the center on the Stiefel manifold,
$\mathcal{PV}(n;\mathcal{H})=\mathcal{V}(n;\mathcal{H})/\mathrm{Z}(n).$ (23)
Thus, the projective Stiefel manifold consists of equivalence classes of $n$-frames in $\mathcal{H}$, where two frames belong to the same class if and only if they differ by a phase factor,
$\bar{V}_{1}=\bar{V}_{2}\iff V_{1}=e^{i\theta}V_{2}\text{ for some }\theta\in\mathbb{R}.$ (24)
We call the equivalence classes projective $n$-frames.
The Stiefel-Grassmann bundle factors through the projective Stiefel-Grassmann bundle, which is the map at the bottom of the commutative diagram
[…]
The projective Stiefel-Grassmann bundle is a principal fiber bundle with symmetry group the projective unitary group $\mathrm{PU}(n)=\mathrm{U}(n)/\mathrm{Z}(n)$. The projective unitary group is a Lie group with Lie algebra $\mathfrak{pu}(n)=\mathfrak{u}(n)/\mathfrak{z}(n)$. [^8] The Lie algebra $\mathfrak{pu}(n)$ is canonically isomorphic to the Lie algebra of trace-free skew-Hermitian $n\times n$ matrices $\mathfrak{su}(n)$, although $\mathrm{PU}(n)$ is not isomorphic to the special unitary group $\mathrm{SU}(n)$. The canonical isomorphism is $\mathfrak{pu}(n)\ni A+\mathfrak{z}(n)\mapsto A-\operatorname{tr}(A)/n\in\mathfrak{su}(n)$, with inverse $\mathfrak{su}(n)\ni A\mapsto A+\mathfrak{z}(n)\in\mathfrak{pu}(n)$.
There is a unique $\mathfrak{pu}(n)$-valued connection $\mathcal{PA}$ on the projective Stiefel manifold making the diagram
${\mathrm{T}_{V}\mathcal{V}(n;\mathcal{H})}$${\mathfrak{u}(n)}$${\mathrm{T}_{\bar{V}}\mathcal{PV}(n;\mathcal{H})}$${\mathfrak{pu}(n)}$$\scriptstyle{\mathcal{A}}$$\scriptstyle{\mathcal{PA}}$
commutative for every $n$-frame $V$. The vertical arrows are the tangent maps of the projections onto the respective orbit spaces. Every tangent vector $Y$ at $\bar{V}$ is the image of a tangent vector $X$ at $V$, and $Y$ is projectively horizontal, that is, $\mathcal{P}\mathcal{A}(Y)=0$, if and only if $\mathcal{A}(X)\in\mathfrak{z}(n)$.
III.C   Projective holonomic gates and projectively parallel transporting Hamiltonians
Let $\mathcal{R}_{t}$ be a closed curve in $\mathcal{G}(n;\mathcal{H})$ at $\mathcal{R}$, and let $\bar{V}$ be any projective frame for $\mathcal{R}$. There exists a unique curve of projective $n$-frames $\bar{V}_{t}$ over $\mathcal{R}_{t}$ that emanates from $\bar{V}$ and whose velocity belongs to the kernel bundle of $\mathcal{PA}$ at all times [21]. This curve is the projective horizontal lift of $\mathcal{R}_{t}$ that starts at $\bar{V}$.
Suppose $V$ is an $n$-frame representing $\bar{V}$. Then $\bar{V}_{t}$ can be lifted to a curve of frames $V_{t}$ for $\mathcal{R}_{t}$ starting at $V$. Any other lift of $\bar{V}_{t}$ to $\mathcal{V}(n;\mathcal{H})$ takes the form $e^{i\theta_{t}}V_{t}$ for some smooth family of phases $\theta_{t}$. That $\bar{V}_{t}$ is projectively horizontal is equivalent to $V_{t}^{\dagger}\dot{V}_{t}\in\mathfrak{z}(n)$. If we write $|v_{j;t}\rangle$ for the $j$th component of $V_{t}$, we can equivalently formulate the projective horizontality condition as
$\langle v_{k;t}|\dot{v}_{l;t}\rangle=i\epsilon_{t}\delta_{kl}\quad(k,l=1,2,\dots,n)$ (25)
for some real numbers $\epsilon_{t}$. Note that $V_{t}$ is conventionally horizontal when $\epsilon_{t}=0$, which we can always achieve by choosing $\theta_{t}$ appropriately. We define the projective holonomy of $\mathcal{R}_{t}$ as the projective unitary operator of $\mathcal{R}$ represented by $V_{\tau}V^{\dagger}$. This definition is independent of both the choice of projective horizontal lift $\bar{V}_{t}$ of $\mathcal{R}_{t}$ and the choice of lift $V_{t}$ of $\bar{V}_{t}$ to $\mathcal{V}(n;\mathcal{H})$. The projective holonomy transforms states with support in $\mathcal{R}$ in the same way as the conventional holonomy.
A Hamiltonian $H_{t}$ with time-translation operator $U_{t}$ is said to be projectively parallel translating if for some (and hence any) $n$-frame $V$ for $\mathcal{R}$, the curve $V_{t}=U_{t}(V)$ satisfies $V_{t}^{\dagger}\dot{V}_{t}\in\mathfrak{z}(n)$ and accordingly represents a projective horizontal curve. Writing $|v_{j}\rangle$ for the $j$th component of $V$, and $|v_{j;t}\rangle$ for $U_{t}|v_{j}\rangle$, the projective parallel transport condition becomes
$\langle v_{k;t}|H_{t}|v_{l;t}\rangle=\epsilon_{t}\delta_{kl}\quad(k,l=1,2,\dots,n)$ (26)
for some real-valued $\epsilon_{t}$; c.f. [9, Eq. (25)]. Unlike the conventional parallel transport condition (3), the projective parallel transport condition is gauge invariant in the sense that if $H_{t}$ is projectively parallel transporting, so is $H_{t}+\epsilon_{t}\mathbbm{1}$ for every real-valued function $\epsilon_{t}$.
III.D   Projective parallel transport in a rotating frame picture
Suppose $V_{t}$ is a curve of $n$-frames. Let $V_{t}^{\textsc{rf}}=R_{t}V_{t}$ be its rotated counterpart, and $A_{t}=i\dot{R}_{t}R_{t}^{\dagger}$ be the rotational potential. Then
$V_{t}^{\dagger}\dot{V}_{t}\in\mathfrak{z}(n)\iff V_{t}^{{\textsc{rf}}\dagger}(d_{t}+iA_{t})V_{t}^{{\textsc{rf}}}\in\mathfrak{z}(n).$ (27)
Thus, in the rotating frame, the projective horizontality condition takes the form
$\langle v_{k;t}^{\textsc{rf}}|d_{t}+iA_{t}|v_{l;t}^{\textsc{rf}}\rangle=i\epsilon_{t}\delta_{kl}\quad(k,l=1,2,\dots,n),$ (28)
while the projective parallel transport condition becomes
$\langle v_{k;t}^{\textsc{rf}}|H_{t}^{\textsc{rf}}-A_{t}|v_{l;t}^{\textsc{rf}}\rangle=\epsilon_{t}\delta_{kl}\quad(k,l=1,2,\dots,n).$ (29)
Here, $\epsilon_{t}$ is an arbitrary smooth real-valued function.
III.E   The isoholonomic inequality for projective holonomic gates
We define the isoholonomic bound of the projective gate $\bar{\Gamma}$ as the minimum of the isoholonomic bounds of the gates that represent $\bar{\Gamma}$,
$L(\bar{\Gamma})=\min\{L(\Gamma):\Gamma\in\bar{\Gamma}\}.$ (30)
Now, suppose $\mathcal{R}_{t}$ is a loop in $\mathcal{G}(n;\mathcal{H})$ with projective holonomy $\bar{\Gamma}$. By the isoholonomic inequality for gates [20],
$\operatorname{Length}[\mathcal{R}_{t}]\geq L(\bar{\Gamma}).$ (31)
Let $\theta_{1},\theta_{2},\dots,\theta_{n}$ be the phases in $[0,2\pi)$ of the eigenvalues of any representative $\Gamma$ of $\bar{\Gamma}$, and set $\theta_{0}=0$. Then,
$L(\bar{\Gamma})=\min_{0\leq k\leq n}\left\{\sqrt{\sum_{j=1}^{n}|\theta_{j}-\theta_{k}|\big{(}2\pi-|\theta_{j}-\theta_{k}|\big{)}}\,\right\}.$ (32)
To see this, first observe that
$L(\bar{\Gamma})=\min_{0\leq\theta<2\pi}\{L(e^{-i\theta}\Gamma)\}.$ (33)
Equation (7) yields
$L(e^{-i\theta}\Gamma)^{2}=\sum_{j=1}^{n}|\theta_{j}-\theta|\big{(}2\pi-|\theta_{j}-\theta|\big{)}.$ (34)
Set $\theta_{n+1}=2\pi$. The function $L(e^{-i\theta}\Gamma)^{2}$, when restricted to $\theta_{k}\leq\theta<\theta_{k+1}$, is a quadratic polynomial with a negative leading coefficient. Such a polynomial never has a value less than the smallest of its limits at the endpoints of the interval. Since $L(e^{-i\theta}\Gamma)^{2}$ is continuous, we conclude that
$L(\bar{\Gamma})^{2}=\min\{L(\Gamma)^{2},L(e^{-i\theta_{1}}\Gamma)^{2},\dots,L(e^{-i\theta_{n}}\Gamma)^{2}\}.$ (35)
Together with Eq. (34), this proves Eq. (32).
III.F   An execution time estimate for projective holonomic quantum gates
For any Hamiltonian driving $\mathcal{R}$ in a loop $\mathcal{R}_{t}$ with conventional holonomy $\Gamma$ and projective holonomy $\bar{\Gamma}$,
$\tau[H_{t};\Gamma]\geq\tau[H_{t};\bar{\Gamma}]=\frac{L(\bar{\Gamma})}{\langle\!\langle\sqrt{I(H_{t};\mathcal{R}_{t})}\,\rangle\!\rangle}.$ (36)
In Ref. [20] we demonstrated that if $\mathcal{R}$ has codimension at least $n$, every gate $\Gamma$ can be generated by a parallel transporting Hamiltonian $H_{t}$ in the time $\tau[H_{t};\Gamma]$. [^9] Whether one can reduce the codimension even more is an open question. Since this holds for any representative $\Gamma$ of $\bar{\Gamma}$ that satisfies $L(\Gamma)=L(\bar{\Gamma})$, we can conclude that if $\mathcal{R}$ has codimension at least $n$, every projective gate $\bar{\Gamma}$ can be generated by a projective parallel transporting Hamiltonian $H_{t}$ in the time $\tau[H_{t};\bar{\Gamma}]$. [^10] This is not as obvious as it may seem because $H_{t}$ and all its shifted versions $H_{t}+\epsilon_{t}\mathbbm{1}$ evolve $\mathcal{R}$ along the same path and with the same speed. Thus, if $H_{t}$ generates $\bar{\Gamma}$ suboptimally, that is, in a time longer than $\tau[H_{t};\bar{\Gamma}]$, the same is true for $H_{t}+\epsilon_{t}\mathbbm{1}$.
IV   Summary
An oftentimes overlooked fact in nonadiabatic holonomic quantum computation is that the parallel transport condition must be covariantly modified to preserve its physical significance when transitioning between different reference frames. In Sec. II, we demonstrated how to properly modify the parallel transport condition when shifting from the laboratory frame to a rotating reference frame.
In Sec. III, we expanded the gauge theory of nonadiabatic holonomic quantum computation into a projective gauge theory, enabling holonomic quantum computation with projective gates. This new framework eliminates the problem of global dynamical phases inherent in conventional nonadiabatic holonomic quantum computation. Additionally, we extended the isoholonomic inequality to projective gates and used this extension to derive a sharp estimate for the execution time of projective holonomic gates. This estimate provides a valuable benchmark for developing faster protocols for holonomic quantum gates.
Acknowledgment
The author would like to thank Niklas Hörnedal for fruitful discussions and constructive comments.